A man is at the top of a tower, and drops a ball. One second later, he throws another ball down with a speed of 20.0 m/s. They hit the ground at the same time. How high (in meters) is the tower?
Answer #1


Assume total time taken by second ball is ť so time for first ball (tt). (t+1) = 27 [fez ist ball) H= 40g (t +12 -H=-200-19+2